A metallic chain 1m long lies on a horizontal surface of a table. The chain starts sliding on the table if 25cm hangs over the edge of a table. The correct value of the coefficient of friction between the table and the chain is1/3 2/3 1/4 1/5

Grade:11

2 Answers

Vikas TU
7 years ago
let m be the mass for 1 m length chain then 25 cm length the mass would be: = > m/4
balacncing the forces we get,
f = mg/4.
u*(3m/4)*g = mg/4
u= 1/3
Akhilesh Patnala
5 years ago
Let us consider [m] as  the mass foe 1m length chain.
Then 25 cm length would have mass ¼th of the total mass . So, mass for 25 cm is m/4 . 
By balancing the forces f=mg/4
we know that f=u[mue]n
u*[3m/4]*g=mg/4.
By solving we get u*=1/3